[Hallc_running] [New Logentry] Thursday RC update

brads at jlab.org brads at jlab.org
Thu Jun 27 12:40:03 EDT 2019

Logentry Text:
There is a "Load shedding" exercise scheduled for 2--3pm this afternoon that will take the machine down in a controlled fashion for at least 1 hour (+ recovery).  We have decided to take advantage of that down time to rotate the SHMS in to 5.7 degrees this afternoon (starting at 2pm).  We'll run the kinematics associated with that SHMS angle during swing and owl, and take another access Friday morning to rotate back out to an angle where the CH can control the spectrometers again.

The earliest possible pass change period is still Saturday, but accelerator downtimes are pushing that towards Sunday -- still TBD though.

---------------  Here's a summary of work since yesterday ---------------

Hall C cryo-target was cooled during Wednesday day shift -- took longer than anticipated due to increased cryo load from Hall A (I gather).  Finished around 4pm.

A power feed to the BSY tripped off around 5pm (no one seems to really understand the root cause yet -- only guesses).  It was reset, but recovery took several hours.

Rocky beam during swing and owl since yesterday afternoon as MCC tries to get acceptable beam to Hall B and continues bringing up Hall A.

Lumi scans on C and LH2 were completed (I believe) -- still need confirmation that we have all the data we want here.

Heep running is still on the todo list for this beam energy (before the pass change).


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3695168
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190627/db0775a4/attachment.html>

More information about the Hallc_running mailing list